In​ 2008, the per capita consumption of soft drinks in Country A was reported to be 19.04 gallons. Assume that the per capita consumption of soft drinks in Country A is approximately normally​ distributed, with a mean of 19.04 gallons and a standard deviation of 5 gallons. Complete parts​ (a) through​ (d) below. a. What is the probability that someone in Country A consumed more than 11 gallons of soft drinks in​ 2008? ​(Round to four decimal places as​ needed.) b. What is the probability that someone in Country A consumed between 2 and 8 gallons of soft drinks in​ 2008? ​(Round to four decimal places as​ needed.) c. What is the probability that someone in Country A consumed less than 8 gallons of soft drinks in​ 2008? ​(Round to four decimal places as​ needed.) d. 98​% of the people in Country A consumed less than how many gallons of soft​ drinks? (2 points) ​(Round to two decimal places as​ needed.
Answer #1

Solution :

Given that ,

mean = \mu = 19.04

standard deviation = \sigma = 5

a) P(x > 11) = 1 - P(x < 11)

= 1 - P[(x - \mu ) / \sigma < (11 - 19.04) /5 )

= 1 - P(z < -1.61)

= 1 - 0.0537

0.9463

Probability = 0.9463

b) P(2 < x < 8) = P[(2 - 19.04)/ )5 < (x - \mu ) /\sigma  < (8 - 19.04) / 5) ]

= P(-3.41 < z < -2.21)

= P(z < -2.21) - P(z < -3.41)

= 0.0136 - 0.0003

0.0133

Probability = 0.0133

c) P(x < 8) = P[(x - \mu ) / \sigma < (8 - 19.04) /5 ]

= P(z < -2.21)

= 0.0136

Probability = 0.0136

d) Using standard normal table ,

P(Z < z) = 98%

P(Z < z) = 0.98

P(Z < 2.05) = 0.98

z = 2.05

Using z-score formula,

x = z * \sigma + \mu

x = 2.05 * 5 + 19.04 = 29.29

98​% of the people in Country A consumed less than 29.29 gallons of soft​ drinks.